Calculate the size of stelar aberration effect in newtonian physics for light coming straight down at speed c and a telescope moving straight sideway at speed v
1
Expert's answer
2020-01-20T05:18:53-0500

According to the rule of addition of velocities in Newtonian physics, but taking into account the independence of the speed of light from the reference frame "sin\theta = \frac{v}{c}"

and since the angle is small, then "\\theta = \\frac{v}{c}.\n\n\u200b"